OpenStudy (anonymous):

For each function find f(-2), f(-0.5), and f(3) Are these correct 8. F(x)= 3/8x -3 1. F= -8 2. F= 4.6 3. F=2.3

OpenStudy (danjs):

f(x) = (3/8) * x - 3 or is that x supposed to be part of that denominator

OpenStudy (danjs):

group the thing with bra-kets for fractions , not sure how you got -8 for the first f(-2), replace x with -2, evaluate for what f(x)=

OpenStudy (anonymous):

I'm sorry I don't understand ): can you show me how to do the first one and I'll do the other two

OpenStudy (danjs):

yes, did i type it in right, the x is multiplied by 3/8 right, it is not 3/(8x) is it

OpenStudy (anonymous):

Correct (: it is multiplied by 3/8

OpenStudy (danjs):

f(-2) is the value of f(x) when x=-2, let x=-2, substitute -2 in for all x's f(-2) = (3/8) *(-2) - 3 = -6/8 - 3

OpenStudy (danjs):

negative six eights minus 3, then simplify more, let 3 become 24/8, to have same denominator , so we can combine fractions

OpenStudy (danjs):

-(6/8) - (24/8) = -30/8

OpenStudy (danjs):

you can reduce that to -15/4 f(-2) = -15/4

OpenStudy (danjs):

----- if x=3 like the last one to figure f(3) you get (9/8) - 3 simplify

OpenStudy (anonymous):

Do I do anything with the -2?

OpenStudy (danjs):

no, -2 is just the x value, f(-2) is the y value for a point on the graph of the function

OpenStudy (anonymous):

My problem is multiplying with fractions

OpenStudy (danjs):

the function f(x) takes any x value you pick along the horizontal x axis, and gives it a y value, it is a rule machine to go from a chosen x value to a corresponding y value

OpenStudy (danjs):

\[\frac{ a }{ b } * \frac{ c }{ d }=\frac{ a*c }{ b*d }\] you just multiply top and bottom seperate like that

OpenStudy (danjs):

so for 9/8 times -2, you can think of -2 as -2/1

OpenStudy (danjs):

the negative 2 multiplies into the numerator

OpenStudy (anonymous):

Oh okay that makes sense (: okay one sec I'm gonna do the problems let me know if they are right (:

OpenStudy (danjs):

to be able to add the fractions, i had to make them both over the same denominator. i multiplied -3 by 8/8 ,

OpenStudy (anonymous):

So I got f(-0.5)= 2.6

OpenStudy (danjs):

let's see what you did . they are mixing decimals and fractions, it is messy you can keep x as -0.5 or make it -1/2

OpenStudy (danjs):

f(-0.5) = (-1.5)/8 - 3

OpenStudy (danjs):

can you do the arithmetic with a calculator? for class

OpenStudy (anonymous):

I did 3/8 multiplied by -1/2 instead of -0.5 then I got -3/16 -3 then I put a 1 under the 3 to make it -3/16 - 3/1 and I got -6/16

OpenStudy (anonymous):

Is it -3/8

OpenStudy (danjs):

you were on the right traack, remember to add or subtract fractions, they have to be both over the same denominator -3/16 - 3*(16/16)

OpenStudy (anonymous):

I don't understand that part ☹

OpenStudy (danjs):

you got the multiplying fractions down, multiply top and bottom together individually addition and subtraction requires the same denominator first

OpenStudy (danjs):

you can add 16ths to 16ths, but you cant add 16ths to 87ths without changing things, the bottom has to be the same

OpenStudy (danjs):

\[\frac{ a }{ c } + \frac{ b }{ c } = \frac{ (a+b) }{ c }\]

OpenStudy (anonymous):

Can we start over Okay so the problem is f(x) = 3/8 (-0.5)-3 F(-0.5) = 3/8 times -0.5 Then what?

OpenStudy (danjs):

we can keep the -1/2 you used, you did that part right (-3/16) - 3 from here you have to put the whole number 3 into 16ths, need same denominator to combine fractions

OpenStudy (anonymous):

Would you be fine doing this with me through snapchat so I can show you what I did

OpenStudy (danjs):

i dont have that

OpenStudy (danjs):

multiply the 3 by (16/16) 16/16 is just 1, you are only multiplying by 1, it doesnt change anything

OpenStudy (danjs):

it becomes 48/16 which is the same as 3

OpenStudy (anonymous):

So -3/16 -3 times 16/16?

OpenStudy (danjs):

yep.. review "common denominators" if you forgot about that ... then you can combine fractions, -3/16 - 48/16

OpenStudy (danjs):

-51/16 or you can approximate with decimals since x is a decimal

OpenStudy (anonymous):

Answer : f(-0.5)= -9

OpenStudy (anonymous):

Oh nevermind you add not multiply woods

OpenStudy (anonymous):

-3.18 is the answer??

OpenStudy (danjs):

-51/16, is that what the calculator says.. looks right

OpenStudy (danjs):

-3.1875

OpenStudy (danjs):

Go for the last one, ill check it at the end, nothing new just a different x value

OpenStudy (anonymous):

Thank you!!!!Can you help me with the last one?

OpenStudy (anonymous):

Ok so f(3) = 3/8 (3)-3 correct?

OpenStudy (anonymous):

6/8 -3?

OpenStudy (anonymous):

Is this the answer?? F(3) = 3.75

OpenStudy (anonymous):

@DanJS

OpenStudy (danjs):

sorry missed this notification, if x is 3 \[f(3) = \frac{ 3 }{ 8 }*3 - 3 ~~~~ =~~~~\frac{ 9 }{ 8 } - \frac{ 24 }{ 8 } \]

OpenStudy (danjs):

= (9-24) /8 = -15/8

OpenStudy (danjs):

you just need to practice adding vs multiplying fractions, this time you added the numerators in the (3/8)*3

OpenStudy (anonymous):

Oh okay thank you so so much!!! Your awesome (:

OpenStudy (danjs):

remember the rules, i think i typed them with the a,b,c,d, general fractions somewhere here

OpenStudy (anonymous):

Would you mind helping me with a new problem?

OpenStudy (anonymous):

And I will thank you(:

OpenStudy (danjs):

ok

OpenStudy (anonymous):

It's fine I'm going to bed thank you for your help! (:

OpenStudy (danjs):

ok, welcome, i may be on tomorrow also if you are working